OpenStudy (lissiewalkin):

HELP PLEASE! An electrochemical cell is made of a left hand compartment consisting of nickel wire in contact with a 1.0M Ni(ClO4)2 solution and a right hand component of copper wire immersed in 1.0M Cu(ClO4)2 1) What is the overall reaction for the electrochemical cell? 2) What is the standard cell potential for this cell? 3) What is the direction of electron flow? (From which electrode to which?) 4) What substance is oxidized? 5) Uf tge [Cu2+] in the right hand compartment is changed to 4.0x10^-3M what is the expected voltage of the electrochemical cell? Can someone please expain

OpenStudy (anonymous):

Ni2+ + 2e →> Ni E= -0.25 Cu2+ + 2e→> Cu E= 0.34 the reduction cell potential indicate that Ni will Oxidize and Cu will reduce. E(not) = E (cathode) - E (anode) E (not) = 0.34 - (-0.25) = 0.59

OpenStudy (anonymous):

E(not) = standard potential

OpenStudy (anonymous):

Electron will flow from Ni (Oxidation occur=loss of electron) to Cu (Reduction occur=gain of electron)

OpenStudy (anonymous):

Ni will oxidized

OpenStudy (anonymous):

1) Ni + Cu^+2 <------> Ni^+2 + Cu

OpenStudy (lissiewalkin):

wow that makes so much more sense now! thanks for the help @waxynaveed! I really appreciate it :]

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

there is potential table which will help u next time to determine which will oxidize and which will reduce..........

OpenStudy (lissiewalkin):

awesome! that table will definitely help. thanks again!

OpenStudy (anonymous):

It's my pleasure :)